# Josephson diodes induced by the loop current states

### v1

@@ Acknowledgements.
I Construction of the Hamiltonian of a Josephson junction
We study a planar Josephson junction (JJ) in which the width of the junction spans unit cells. The two superconducting regions each have a length of unit cells, while the central non-superconducting region extends over unit cells as shown in Fig. S1. Thus the Hamiltonian of the JJ can be expressed as . Here, and describe the two superconducting regions on the left and right sides, which consists of the model with the loop current state described in the main text and a local -wave superconducting pairing with a phase bias . describes the normal region in the middle with the superconducting pairing term set to 0. and correspond to the coupling between the normal region and the left and right superconducting region. Moreover, a staggered lattice potential can be described by an extra term and an out-of-plane electric field for the bilayer system can be described by a term . The periodic boundary condition is adopted along the width of the junction and the total Hamiltonian is diagonalized numerically. Then the Josephson currents can be calculated as
@@ I Construction of the Hamiltonian of a Josephson junction
II CPR for the JJ constructed by the monolayer HM model
For the JJ constructed by the monolayer HM model, the CPR shows no diode effect since the HM model has the inversion symmetry ( for the monolayer system) as shown in Fig. S2.
III CPR for the armchair JJ constructed by the bilayer models
As mentioned in the main text, the JDE can be achieved by applying an out-of-plane electric field to the bilayer HM with the same loop current direction or stacking the two layers of the HM with opposite loop current directions in the zigzag JJs. However, if the JJs are constructed along the armchair edge, the JDE for both cases disappears, which is related to the mirror symmetry of the armchair JJs. Here, we provide the numerical evidence to verify these results in Fig. S3, where the Josephson current of the armchair JJ is an odd function of the phase bias and thus does not show any diode effect.
IV Symmetry analysis
In this section, we demonstrate the reason why the JDE realized in the zigzag JJs disappears in the armchair JJs. From the symmetry perspective, if the system remains invariant under a certain operation that simultaneously changes the sign of the current operator, it implies that the critical current must have the same magnitude in opposite directions, resulting in the absence of the JDE. We find that such symmetry operations can be constructed by the following three operations: the mirror reflection (), the time-reversal () and inversion ( in monolayer model). Both the and operations can change the current direction. represents the mirror reflection about the plane perpendicular to the -axis, and the definition of the directions is shown in Fig. S4, where corresponds to the zigzag direction and represents the armchair direction. For the zigzag JJ, changes the current direction while does not. Conversely, for the armchair model, changes the current direction, but does not.
For JJs constructed by different models with difference junction directions, the possible symmetry operations that prevent the emergence of the JDE are summarized in Table. 1. For example, for the armchair JJ constructed by the HM model with staggered lattice potential, operations , and all change the direction of the current. Therefore, the combined operation changes the direction of the current while keeping the system invariant. Hence, the JDE is absent in the armchair JJ constructed by the HM model with staggered lattice potential. For zigzag JJ constructed by the same model, we can not find any symmetry operations that also changes the direction of the current which is represented by the ‘’ sign in the table.
monolayer | operations | JDE |
armchair HM | No | |
armchair MHM | No | |
zigzag HM | No | |
zigzag MHM | Yes | |
monolayer with staggered potential | operations | JDE |
armchair HM | No | |
armchair MHM | No | |
zigzag HM | Yes | |
zigzag MHM | Yes | |
bilayer HM with same current direction and interlayer potential | operations | JDE |
armchair HM | No | |
zigzag HM | Yes | |
bilayer HM with opposite current direction | operations | JDE |
armchair HM | No | |
zigzag HM | Yes |

### v3

@@ Acknowledgements.
IX Construction of the Hamiltonian of a Josephson junction
We study a planar Josephson junction (JJ) in which the width of the junction spans unit cells. The two superconducting regions each have a length of unit cells, while the central non-superconducting region extends over unit cells as shown in Fig.Β S1. Thus the Hamiltonian of the JJ can be expressed as
where represents the Hamiltonian of monolayer () or bilayer () system. For the monolayer system, it can be written as
describing the monolayer Haldane or modified Haldane models. For the bilayer system, it can be expressed as
describing the bilayer Haldane or modified Haldane model coupled through the interlayer coupling . Thus, and correspond to the Hamiltonian of the superconducting region on the left and right sides. describes the normal region in the middle with the superconducting pairing term set to 0. and correspond to the spin-conserving hoppings between the superconducting and normal sections which include both the nearest-neighbor and next-nearest-neighbor hoppings and take the same values and patterns as those in the normal region. The superconducting phase bias for the JJ is assigned to the left superconducting region as so that the phase in the right superconducting region is set to . The amplitude of the order parameter is taken as for all calculations in this paper. represents the chemical potential of the entire material, which is set to -0.4t in the bilayer model, and set to 0 in the monolayer system. Moreover, a staggered lattice potential can be described by an extra term added into the monolayer Hamiltonian and an out-of-plane electric field for the bilayer system can be described by a term . The periodic boundary condition is adopted along the width of the junction and the total Hamiltonian is diagonalized numerically. Then the Josephson currents can be calculated as
@@ IX Construction of the Hamiltonian of a Josephson junction
X CPR for the JJ constructed by the monolayer HM model
For the JJ constructed by the monolayer HM model, the CPR shows no diode effect since the HM model has the inversion symmetry ( for the monolayer system) as shown in Fig.Β S2.
XI CPR for the armchair JJ constructed by the bilayer models
As mentioned in the main text, the JDE can be achieved by applying an out-of-plane electric field to the bilayer HMs with the same loop current direction or stacking the two layers of the HMs with opposite loop current directions in the zigzag JJs. However, if the JJs are constructed along the armchair edge, the JDE for both cases disappears, which is related to the mirror symmetry of the armchair JJs. Here, we provide the numerical evidence to verify these results in Fig.Β S3, where the Josephson current of the armchair JJ is an odd function of the phase bias and thus does not show any diode effect.
XII The effect of the chemical potential and pairing strength
To explore the effect of the chemical potential and the pairing amplitude to the diode efficiency for the JJs constructed by the bilayer HM, we calculate for various values of and . The pairing amplitude is varied from to and the chemical potential is varied from to in the bilayer model with other parameters unchanged. The results are shown in Fig.Β S4 where the parameters used in the paper are marked with black dots. We can see that the diode efficiency increases significantly when and weak , while it remains relatively low in other parameter regions.